OpenStudy (anonymous):

Solve. 7^x – 5 = 7^4x + 7

OpenStudy (anonymous):

x = –4 x = negative 12 over 5 x = 2 over 3 x = 4

OpenStudy (anonymous):

Is it \(\large 7^{x - 5} = 7^{4x + 7} ?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

They have common bases, so just drop the bases. x - 5 = 4x + 7 Can you solve now?

OpenStudy (anonymous):

is it 2/3?

OpenStudy (anonymous):

or just 4? I' don't think I'm doing the math right

OpenStudy (anonymous):

4x + 7 = x - 5 3x = -12 x = -4